OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. (Do not type the variable in the blank.) 4x2 + ___x + 49

OpenStudy (anonymous):

is this 14?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

its 14

OpenStudy (anonymous):

the polynominal is (2x+7)^2

OpenStudy (anonymous):

oops its 28

OpenStudy (anonymous):

wait i thought it's be 14 because half is 7 & 7 squared is 49?

OpenStudy (anonymous):

oldnt the answer be 14?

OpenStudy (anonymous):

wouldnt *

OpenStudy (anonymous):

alright all this switching is causingme problems

OpenStudy (anonymous):

im prettty sure its 14 haha

OpenStudy (anonymous):

\[(a+b)^2=a^2+2ab+b^2\] if a=2x \[a^2=4x^2\] if b = 7 \[b^2=49\] so these are your a and b's \[a^2+2ab+b^2=4x^2+2(2x)(7)+49\] \[4x^2+28x+49\]

OpenStudy (anonymous):

so its 28?

OpenStudy (lgbasallote):

it's not 14..you need to divide everything by 4 before you do completing the square

OpenStudy (anonymous):

this is so confusing.

OpenStudy (lgbasallote):

i know right

OpenStudy (anonymous):

without the 4, it'd be 14

OpenStudy (anonymous):

well it doesn't help when someone keeps chiming in o wait this answer this answer.... So i had to start over from scratch

OpenStudy (anonymous):

haha yeah

OpenStudy (lgbasallote):

here's a helpful hint the general equation is \(ax^2 + bx + c\) \[(\frac{b}{2})^2 = c\] since we're looking for b we do formula transformation \[\frac b2 = \sqrt c\] \[b = 2\sqrt c\] now our c would be 49/4 because we divided by 4 in the start \[b = 2 \sqrt{49/4}\] \[b = 2(7)/2)\] \[b = 7\] could it be @Outkast3r09 made a mistake or did i...